OpenStudy (anonymous):

Solve the proportion below. x/12.6= 45/63 x=

OpenStudy (anonymous):

If it is x/12.6 = 45/63, then x=(45)(12.6)/63=9 If it is (x/12)6 = 45/63, then x=(45)(2)/63=1.43 If it is x/(12)(6) = 45/63, then x=(45)(12)(6)/63=51.43

OpenStudy (anonymous):

x=9